IN 20 gm ring here is 60 % gold and 40 % silver then how much gold is mixed such that gold in ring becomes 80 %
Answer by flame8855(424)

in 20 gm ring
0.6 X 20 = 12 g gold
o.4X20 = 8g silver
when we add x number of gold
total weight is = 20 + x
total number of gold = x+12
80 % gold
0.8= number of gold divided by total mass
0.8 = (x+12)/(20+x)
x= 20
so we need to ad 20 g of gold